Shift report 20030304 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Mar 04 2003 - 07:57:21 EST

  • Next message: Flemming Videbaek: "Nuclear Physics Gordon Conference"
    ------------------------------------------------------------
    Date: Tue Mar  4 07:57:20 2003, Shift: 00:00-08:00
    Supervisor: FR, with: 
    Log book#14 pages 45 - 45
    
    Shift Summary:
    ------------------------------------------------------------
    
    1) Beam was on at the beginning of the shift but background was too high (TOFW/ZDC~100). Voted with PHENIX to end the store. The beam was dumped at 1:30. Then MCR announced a foil change in the tandem. After this, there were a few
    unsuccessful attempts at ramping. Since about 6:30 "RHIC Quench Link Interlock" problems. Current status: Machine setup.
    
    2) Runs 8196-8198 are TPC calibration runs.
    
    3) No detector problems noticed.
    
    Run plan: FS at 4deg 1/2-B field, MRS at 40deg 1050-A field.
    The goal for this setting is to collect 100k trig 2 events.
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Tue Mar 04 2003 - 07:58:32 EST